OpenStudy (anonymous):

find the two number whose sum is 10 abd the sum of whose square is minimum

OpenStudy (dumbcow):

x+y = 10 S = x^2 + y^2 to minimize S get in terms of x, differentiate and set equal to 0 solve for x y= 10-x S= x^2 + (10-x)^2 go from there